# Few-Shot Learning on Graphs: from Meta-learning to Pre-training and Prompting

### v3

@@ 3.3 Parameter-efficient downstream adaptation
Parameter-efficient fine-tuning (PEFT). These methods only tune a selected subset of parameters from the original pre-trained model or from newly added modules. Hence, the number of parameters requiring updating is significantly reduced, which proves more efficient and feasible for few-shot learning. Prominent PEFT techniques include adapter tuning Houlsby et al. (2019) and Low-Rank Adaptation (LoRA) Hu et al. (2021). Adapter tuning Houlsby et al. (2019) inserts small neural network modules called adapters into certain layers of a pre-trained model, while LoRA Hu et al. (2021) utilizes low-rank matrices to approximate parameter updates. In both techniques, only the parameters of the new modules are updated during fine-tuning, while the original pre-trained weights remain frozen.
@@
For pre-trained graph models, AdapterGNN Li et al. (2023a) introduces adapter-based fine-tuning into graph neural networks. This is achieved via two parallel adapters that process inputs before and after message passing. In a similar spirit, G-Adapter et al.Gui et al. (2023) refines the graph transformer architecture with adapters, and further incorporates structural information into the adapter. In a different graph-assisted application, Liu et al. Liu et al. (2023d) focus on the molecule-to-text generation task and propose molecular graph language modeling. They design a cross-modal projector aimed at bridging the gap between graph structural and textual representations, while integrating a uni-modal LoRA module into their language model.
3.4 Discussion
@@ 4 Hybrid Approaches
As depicted in Fig. 4(a), hybrid approaches adopt typical contrastive or generative pretext tasks to pre-train a graph encoder. Subsequently, the pre-trained model is adapted in conjunction with meta-learning, as illustrated in Fig. 4(b,c). Specifically, VNT Tan et al. (2023) employs a pre-trained model to reduce reliance on the base classes. It further proposes structure prompts, serving as virtual nodes to tailor the pre-trained node embeddings for meta-training tasks, further leveraging the insights from the base classes. However, VNT is specifically designed for node-level tasks only, falling short of handling edge- or graph-level tasks. In contrast, ProG Sun et al. (2023) reformulate the downstream node- and edge-level tasks as graph-level tasks, and proposes prompt graphs with specific nodes and structures to guide different tasks. Similar to VNT, it incorporates meta-learning to acquire initial prompts for handling different tasks. On text-attributed graphs, G2P2* Wen and Fang (2023b) extends G2P2 Wen and Fang (2023a) with conditional prompt tuning, which employs a meta-net to generate a distinct prompt vector for each node. The meta-net is trained using labeled data from base classes, facilitating knowledge transfer to novel classes.
@@ 5 Future Directions
5.1 Future avenues in problem setting
Few-shot learning on large-scale graphs. Previous works predominantly concentrate on small graph structures such as molecule graphs Hu et al. (2020a), which consist of tens or hundreds of nodes only. However, many real-world applications deal with much larger graph structures, such as social networks, interaction network for recommender systems, and protein or gene networks. Challenges on these larger graphs encompass not only the immediate issues of computational efficiency and scalability, but also more finer-grained adaptation strategies to deal with potential variations among distant localities on a larger graph.
Few-shot learning on complex graphs. Most of the current works focus on simple homogeneous structures, commonly exemplified by 2D graphs, uni-modal graphs, and static graphs. However, the practical applicability of few-shot learning extends beyond these limitations to include more complex structures, such as 3D graphs Liu et al. (2021a), multi-modal graphs Li et al. (2023b), and dynamic graphs Yang et al. (2022), each with important applications in scientific, biomedical or social domains. This expansion in scope highlights the need for novel methodologies and techniques to effectively address the challenges posed by each of these intricate graph structures.
Few-shot learning on cross-domain graphs. Cross-domain few-shot learning involves transferring knowledge from one domain to another, emphasizing the generalization across diverse scenarios Yao et al. (2020). Graphs from different domains may vary in size, topology and semantic interpretation, which may call for specialized domain adaptation techniques to mitigate domain shift.
5.2 Future avenues in techniques
@@
Foundation models on graphs. Foundation models pre-trained on large-scale data in NLP have shown powerful performance across various kinds of tasks in diverse application domains. Similarly, a pre-trained graph foundation model holds significant promise for diverse few-shot learning tasks on graphs. Toward graph foundation models, we aim to improve the transferability of current pre-training efforts on graphs, enabling their application in diverse domains as well as different task types. On one hand, improving transferability across domains requires not only mitigating the issue of domain shift on cross-domain graphs (see Sect. 5.1), but also bridging the gap created by heterogeneity in graph structures. This heterogeneity stems from different kinds of complex structures, including 3D, multi-modal and dynamic graphs (see Sect. 5.1). Unifying these structures is critical to the universal applicability of graph foundation models; in particular, multi-modal graphs pave the way for integrating graph foundation models with language- and vision-based models. On the other hand, there is a pressing need to enhance transferability across task types, extending beyond the conventional tasks of link prediction, node classification, and graph classification. More variety of task types, such as regression, graph editing and generation Liu et al. (2023b), can be crucial to material engineering and drug development.
6 Conclusions
In this survey, we offer a systematic yet concise review of the literature on few-shot learning on graphs. Based on predominant techniques, we categorize these studies into meta-learning approaches, pre-training approaches and hybrid approaches. For each category, we present representative works and discuss their respective advantages and drawbacks. Finally, we outline promising future directions, aiming to stimulate further exploration in this rapidly evolving field.

### v4

@@ V-C2 Parameter-Efficient Fine-Tuning (PEFT)
These methods only tune a selected subset of parameters from the original pre-trained model or from newly added modules. Hence, the number of parameters requiring updating is significantly reduced, which proves more efficient and feasible for few-shot learning. Prominent PEFT techniques include adapter tuning [188] and Low-Rank Adaptation (LoRA) [189]. Adapter tuning [188] inserts small neural network modules called adapters into certain layers of a pre-trained model, while LoRA [189] utilizes low-rank matrices to approximate parameter updates. In both techniques, only the parameters of the new modules are updated during fine-tuning, while the original pre-trained weights remain frozen.
@@
For pre-trained graph models, AdapterGNN [46] introduces adapter-based fine-tuning into graph neural networks. This is achieved via two parallel adapters that process inputs before and after message passing. In a similar spirit, G-Adapter [47] refines the graph transformer architecture with adapters, and further incorporates structural information into the adapter. In a different graph-assisted application, MolCa [81] focus on the molecule-to-text generation task and propose molecular graph language modeling. They design a cross-modal projector aimed at bridging the gap between graph structural and textual representations, while integrating a uni-modal LoRA module into their language model.
V-D Discussion
@@ VI Hybrid Approaches
As depicted in Fig. 7(a), hybrid approaches adopt typical contrastive or generative pretext tasks to pre-train a graph encoder. Subsequently, the pre-trained model is adapted in conjunction with meta-learning, as illustrated in Fig. 7(b,c). Specifically, VNT [87] employs a pre-trained model to reduce reliance on the base classes. It further proposes structure prompts, serving as virtual nodes to tailor the pre-trained node embeddings for meta-training tasks, further leveraging the insights from the base classes. However, VNT is specifically designed for node-level tasks only, falling short of handling edge- or graph-level tasks. In contrast, ProG [49] reformulate the downstream node- and edge-level tasks as graph-level tasks, and proposes prompt graphs with specific nodes and structures to guide different tasks. Similar to VNT, it incorporates meta-learning to acquire initial prompts for handling different tasks. On text-attributed graphs, G2P2* [82] extends G2P2 [44] with conditional prompt tuning, which employs a meta-net to generate a distinct prompt vector for each node. The meta-net is trained using labeled data from base classes, facilitating knowledge transfer to novel classes.
@@ VII Future Directions
VII-A Future Avenues in Problem Settings
Structure scarcity learning on graphs. As elaborated in Sect. III-B, the issue of structure scarcity in graphs is prevalent across various real-world scenarios. Existing solutions [132, 136] to this challenge typically rely on meta-learning approaches, which learns a prior from tasks involving structurally rich instances. However, meta-learning often assumes that base and novel tasks are independent and identically distributed (i.i.d.). When dealing with structure scarcity, the base tasks are constructed from structure-rich instances, while the novels tasks are associated with structure-scarce instances. The discrepancy may lead to distinct distributions between them, violating the i.i.d. assumption. One potential solution is to leverage pre-training to learn priors from the base tasks. Since downstream tasks typically differ from those used in pre-training, specialized adaptation techniques are often introduced to bridge the gap between them, rather than making the i.i.d. assumption.
Few-shot learning on large-scale graphs. Many previous works [40, 19] are applied to small-scale graph structures such as molecular graphs [190], which consist of tens or hundreds of nodes only. However, many real-world applications deal with much larger graph structures, such as social networks, e-commerce systems, and protein or gene networks. Though some existing works [191, 192] reduce the computational cost to a certain extent, it remains a significant issue on very large graphs. Additionally, challenges persist in developing finer-grained adaptation strategies to deal with potential variations among distant localities on a large graph.
Few-shot learning on complex graphs. Most of the current works focus on simple homogeneous graphs, commonly exemplified by 2D graphs, uni-modal graphs, and static graphs. However, the practical applicability of few-shot learning extends beyond these limitations to include more complex structures, such as 3D graphs [193], multi-modal graphs [194], and dynamic graphs [137], each with important applications in scientific, biomedical or social domains. This expansion in scope highlights the need for novel methodologies and techniques to effectively address the challenges posed by these complex graph structures. However, existing methods typically focus on a single aspect of these complexities or are tailored to specific applications [193, 111], lacking a unified approach that integrates various aspects and general applications.
Few-shot learning on cross-domain graphs. Graphs are ubiquitous across various domains, from social networks and commerce networks to molecular graphs and protein-protein interaction networks. Each domain’s graphs exhibit unique semantic meanings and topological structures. For example, in a social network, a node’s features may represent a user’s profile, while the links denote the connections between users. In contrast, in a molecular graph, a node’s features represent the attributes of an atom, and the edges indicate chemical bonds. This diversity highlights the need for cross-domain few-shot learning methods. Some current solutions rely on textual descriptions on graphs to bridge different domains via LLMs [45], but this creates a heavy dependency on LLMs and does not work for text-free graphs [181]. Other approaches ignore the topological differences across various domains [195, 181], calling for specialized domain adaptation techniques to mitigate both feature and structure variations across domains.
VII-B Future Avenues in Techniques
@@
Foundation models on graphs. Foundation models pre-trained on large-scale data in NLP have shown powerful performance across various kinds of tasks in diverse application domains. Similarly, a pre-trained graph foundation model holds significant promise for diverse few-shot learning tasks on graphs. Toward graph foundation models, we aim to improve the transferability of current pre-training efforts on graphs, enabling their application in diverse domains as well as different task types. On one hand, improving transferability across domains requires not only mitigating the issue of domain shift on cross-domain graphs, but also bridging the gap created by heterogeneity in graph structures as discussed in Sect. VII-A. This heterogeneity stems from different kinds of complex structures, including 3D, multi-modal and dynamic graphs. Unifying these structures is critical to the universal applicability of graph foundation models; in particular, multi-modal graphs pave the way for integrating graph foundation models with language- and vision-based models. On the other hand, there is a pressing need to enhance transferability across task types, extending beyond the conventional tasks of link prediction, node classification, and graph classification. More variety of task types, such as regression, graph editing and generation [196], can be crucial to material engineering and drug development.
VIII Conclusions
In this survey, we offer a systematic review of the literature on few-shot learning on graphs. Specifically, we develop two comprehensive taxonomies for few-shot learning on graphs, classifying existing studies according to their problems and techniques. Based on the problems, we divide these studies into two categories: label-scarcity problem and structure-scarcity problem. For each category, we summarize their problem definitions and their respective applications. Based on the techniques, we categorize the literature into meta-learning, pre-training, and hybrid approaches. For each category, we synthesize representative works and discuss their advantages and disadvantages. Finally, we outline promising future directions, aiming to stimulate further exploration in this rapidly evolving field.